OpenStudy (anonymous):

Ok I have two questions. Can work one at a time. One zero of f(x)=x^3+2x^2-5x-6 is 2. What are the other zeros of the function? a.) 1 and 3 b.) -1 and -3 c.) 2 and -3 d.) 2 and 3 Which of the following quartic functions has x=-2 and x=-3 as its only two real zeros? a.) x^4-5x^3+7x^2-5x+6 b.) x^4+5x^3+7x^2+5x+6 c.) x^4+5x^3+7x^2-5x-6 d.) x^4-5x^3+7x^+5x-6

OpenStudy (michele_laino):

The complete factor of f(x), is: \[f(x)=(x-2)*(x+1)*(x+3)\] so, you have other zeroes, try to write them, please

OpenStudy (anonymous):

1 and 3

OpenStudy (michele_laino):

Sorry, incorrect answer!

OpenStudy (michele_laino):

please , you have to solve two linear equation, namely: \[x+1=0,x+3=0\]

OpenStudy (michele_laino):

re-try, please!

OpenStudy (anonymous):

x = -1 and x = -3

OpenStudy (michele_laino):

That's right!

OpenStudy (michele_laino):

Ok!, For your second answer, you have to substitute your number, namely -2 and -3 in all the equations indicated. When you get zero after both substitutions, you have found the right equation. Please try!